# Optimizing Parameter Estimation for Electrochemical Battery Model: A Comparative Analysis of Operating Profiles on Computational Efficiency and Accuracy

### v2

@@ 3.5 L4: Second Smallest Value Obtained for Each Column
It is important to notice that the minima is not the only important part but also how far the next smallest values are. Indeed, if the next small error value is relatively close to the minimum one, it might still be a reasonable error mark but the parameter error and total time requirements might be more important for this sub-optimal solution compared to the optimal one. For Scenario 1 (column 1 in Fig. 6), for instance, the next error mark to the minimum appears for Case 21 with an 8% of error difference. However, while Case 1 (the minimum) requires the use of all the available datasets, i.e. {C/5,C/2,1C,P,DST}, Case 21 resorts to dataset {C/2,1C} at the price of 8% more of output error. This is exemplified in the last 3 columns of Table 6 where the minimal dataset for Scenario 2 corresponds to the 2nd minimum (row 2, column 2), for which the corresponding Case is 21 (row 2, column 3) with associated 8% error (row 2, column 4). A similar reasoning can be used for the other Scenarios. In most situations, the required data will be reduced by 1 to 3 datasets since all situations above Case 16 (see Figure 6) require the full 5 datasets (Case 1), 4 datasets (Cases 2 to 6) or 3 datasets (Cases 7 to 16) and data reduction can go up to 2 datasets (Cases 17 to 26) or 1 dataset (Cases 27 to 31) if an error within the 5 smallest values is acceptable (below 15% of error). An exception is Scenario 11, where the minimum error is found for Case 11 and the shortest dataset that can be obtained within the first 5 smallest errors is for Case 12, with the 5th smallest error of 16% and still keeping 3 datasets (from {C/5,1C,DST} to {C/5,P,DST}). In general, dataset reduction is dependent on the original dataset length, where longer datasets are more prone to be reduced with small errors than shorter datasets. In the limit where only 1 dataset is used (Cases 27 to 31), no further reduction is possible and in some Cases, the corresponding error is above the 5 smallest errors for that column (for Scenarios 28 and 30 for instance). These notions are naturally exploited below when a cost function is used to find the best combination of conditions to optimize a given set of metrics.
@@ 3.7 L6: Parameter Errors
Table 7 shows the component-wise relative error of each parameter component in the parameter vector Equation (12) as defined in Equation (16) shown in column 2 to 10, where (with ) in Table 7 just specifies the considered parameter component, the average parameter error in column 11, the Euclidean distance of the parameter vector errors (as in Equation (17)) in column 12 . The important notion is the distance between the identified parameters and their nominal values. Starting from the component-wise errors, we see that minimum errors appear for the parameters {} in the Cases {19, 28, 28, 25, 26, 4, 3, 19, 21}, respectively, with errors of {1.44, 0.47, 0.08, 0.24, 2.94, 0.74, 0.70, 0.07, 0.88}%, respectively. Thus, the maximum error in a parameter component is around 3% for . Most component minimum errors show up for datasets with 1 or 2 profiles (Cases 17 to 31) except for the errors for and for Cases 4 and 3, respectively. The minimum of the average of these errors for each Case (column 11 in Table 7) corresponds to 8.45% for Case 4. Even if this gives an idea of the level of error of each parameter component, we also look at the Euclidean distance between the estimated parameter vector and the nominal one (column 12 in the table).
@@ 3.8 L7: Overall Best Datasets for Given Performance Metrics
The optimal operating condition for each conditions is summarized in Table 9. From the results in Table 9, it can be observed that if we only consider minimizing the output voltage error (O1), Case 1 {C/5,C/2,1C,P,DST} is the optimal choice as it encompasses all operating conditions. However, it also has the longest computation time (55 hours). Interestingly, although the model voltage output error for Case 1 is the smallest, its parameter error is not the lowest. This highlights the significant challenges in electrochemical model parameter estimation and the issue of over-parameterization, where multiple combinations of parameters can result in similarly accurate model outputs. In other words, large parameter errors may still yield small model voltage output errors.
@@
Based on the above results, the optimal operating condition can be selected according to specific needs. For example, if the primary concern is model voltage output error and computation time (O5), Case 21 {C/2,1C} is the recommended option, as this aligns with the most common practical requirements. Due to the issue of over-parameterization, practical applications often prioritize minimizing model voltage output error over parameter accuracy. However, if more accurate battery parameter information is needed, such as for explaining internal resistance increases caused by aging, Case 4{C/5,C/2,P,DST} should be considered despite the increased computation time. Researchers can select the most suitable operating condition based on their specific requirements.
@@ 3.9 Limitation
This study focuses exclusively on lithium-ion batteries utilizing NMC materials, and the conclusions drawn are specifically applicable to these systems. Although the comprehensive analytical methodology presented here has been successfully demonstrated for NMC-based batteries, it is equally applicable to other battery chemistries, including Lithium Iron Phosphate (LFP) and other batteries. By applying the same approach to LFP and other type batteries, researchers may streamline testing protocols and obtain corresponding performance insights.

### v3

@@ 3.5 L4: Second Smallest Value Obtained for Each Column
It is important to notice that the minima is not the only important part but also how far the next smallest values are. Indeed, if the next small error value is relatively close to the minimum one, it might still be a reasonable error mark but the parameter error and total time requirements might be more important for this sub-optimal solution compared to the optimal one. For Scenario 1 (column 1 in Figure 6), for instance, the next error mark to the minimum appears for Case 21 with an 8% of error difference. However, while Case 1 (the minimum) requires the use of all the available datasets, i.e. {C/5,C/2,1C,P,DST}, Case 21 resorts to dataset {C/2,1C} at the price of 8% more of output error. This is exemplified in the last 3 columns of Table 6 where the minimal dataset for Scenario 2 corresponds to the 2nd minimum (row 2, column 2), for which the corresponding Case is 21 (row 2, column 3) with associated 8% error (row 2, column 4). A similar reasoning can be used for the other Scenarios. In most situations, the required data will be reduced by 1 to 3 datasets since all situations above Case 16 (see Figure 6) require the full 5 datasets (Case 1), 4 datasets (Cases 2 to 6) or 3 datasets (Cases 7 to 16) and data reduction can go up to 2 datasets (Cases 17 to 26) or 1 dataset (Cases 27 to 31) if an error within the 5 smallest values is acceptable (below 15% of error). An exception is Scenario 11, where the minimum error is found for Case 11 and the shortest dataset that can be obtained within the first 5 smallest errors is for Case 12, with the 5th smallest error of 16% and still keeping 3 datasets (from {C/5,1C,DST} to {C/5,P,DST}). In general, dataset reduction is dependent on the original dataset length, where longer datasets are more prone to be reduced with small errors than shorter datasets. In the limit where only 1 dataset is used (Cases 27 to 31), no further reduction is possible and in some Cases, the corresponding error is above the 5 smallest errors for that column (for Scenarios 28 and 30 for instance). These notions are naturally exploited below when a cost function is used to find the best combination of conditions to optimize a given set of metrics.
@@ 3.7 L6: Parameter Errors
Table 7 shows the component-wise relative error of each parameter component in the parameter vector Eq. (12) as defined in Eq. (16) shown in column 2 to 10, where (with ) in Table 7 just specifies the considered parameter component, the average parameter error in column 11, the Euclidean distance of the parameter vector errors (as in Eq. (17)) in column 12 . The important notion is the distance between the identified parameters and their nominal values. Starting from the component-wise errors, we see that minimum errors appear for the parameters {} in the Cases {19, 28, 28, 25, 26, 4, 3, 19, 21}, respectively, with errors of {1.44, 0.47, 0.08, 0.24, 2.94, 0.74, 0.70, 0.07, 0.88}%, respectively. Thus, the maximum error in a parameter component is around 3% for . Most component minimum errors show up for datasets with 1 or 2 profiles (Cases 17 to 31) except for the errors for and for Cases 4 and 3, respectively. The minimum of the average of these errors for each Case (column 11 in Table 7) corresponds to 8.45% for Case 4. Even if this gives an idea of the level of error of each parameter component, we also look at the Euclidean distance between the estimated parameter vector and the nominal one (column 12 in the table).
@@ 3.8 L7: Overall Best Datasets for Given Performance Metrics
The optimal operating condition for each conditions is summarized in Table 9. From the results in Table 9, it can be observed that if we only consider minimizing the output voltage error (O1), Case 1 {C/5,C/2,1C,P,DST} is the optimal choice as it encompasses all operating conditions. However, it also has the longest computation time (55 hours). Interestingly, although the model voltage output error for Case 1 is the smallest, its parameter error is not the lowest. This highlights the significant challenges in electrochemical model parameter estimation and the issue of over-parameterization, where multiple combinations of parameters can result in similarly accurate model outputs. In other words, large parameter errors may still yield small model voltage output errors. This discrepancy may also be attributed to differences in parameter sensitivity, as some parameters exhibit inherently low sensitivity and are therefore difficult to estimate accurately. In addition, the feasible range of battery parameters can further influence the parameter estimation results.
@@
Based on the above results, the optimal operating condition can be selected according to specific needs. For example, if the primary concern is model voltage output error and computation time (O5), Case 21 {C/2,1C} is the recommended option, as this aligns with the most common practical requirements. In many real-world scenarios, the exact range of parameter values, potential over-parameterization, and parameter sensitivity are of secondary importance; the primary focus is instead on reducing the model’s voltage error. However, if more accurate battery parameter information is needed, such as for explaining internal resistance increases caused by aging, Case 4{C/5,C/2,P,DST} should be considered despite the increased computation time. Researchers can select the most suitable operating condition based on their specific requirements.
In fact, the findings of this study can provide direct guidance for such applications. For instance, in model-based SOC estimation methods, the SOC accuracy is strongly influenced by the model’s voltage prediction error. Therefore, the optimal current combinations identified for minimizing voltage error (e.g., O1 and O5 in Table 9) are particularly well-suited for SOC estimation tasks. Conversely, in the context of SOH re-calibration, capacity-related parameters play a more critical role. Hence, current combinations that yield lower errors in capacity-related parameters (e.g., O2 and O6 in Table 9) are preferable for SOH estimation. Accordingly, our systematic evaluation of 31 operating condition combinations offers practical insights for selecting appropriate test profiles tailored to different application goals (SOC vs. SOH).
@@ 3.9 Limitation
This study focuses exclusively on lithium-ion batteries utilizing NMC materials, and the conclusions drawn are specifically applicable to these systems. Although the comprehensive analytical methodology presented here has been successfully demonstrated for NMC-based batteries, it is equally applicable to other battery chemistries, including Lithium Iron Phosphate (LFP) and other batteries. By applying the same approach to LFP and other type batteries, researchers may streamline testing protocols and obtain corresponding performance insights. The number of PSO iterations may also have a potential impact on the results. In this study, we adopted a commonly used setting of 300 iterations for the analysis. Using a larger number of iterations could reduce the error for certain operating conditions, but it would also lead to a longer computational time.
@@ 4 Conclusions
Acknowledgments
This work was supported by the Research Foundation - Flanders (FWO) (grant numbers 1252326N, 2025).
Author contribution
Feng Guo: Conceptualization, Methodology, Software, Formal analysis, Investigation, Visualization, Writing – original draft. Luis D. Couto: Conceptualization, Methodology, Visualization, Writing – original draft, Writing – review & editing. Khiem Trad: Conceptualization, Writing – review & editing. Grietus Mulder: Writing – review & editing. Keivan Haghverdi: Writing – review & editing. Guillaume Thenaisie: Writing – review & editing.
Data Availability
The implementation of the models used in this paper can be accessed at https://github.com/FrankSuperG/CPG-SPMT.